Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2016-11-20

Observers: K. Takáts, Ósmar Rodríguez
Support team: Ken Smith
TAT team: Joe Anderson, Kate Maguire

Daytime calibrations:
SOFI:
PESSTO_SOFI_specflat
PESSTO_SOFI_specArc
PESSTO_SOFI_FlatJHK

EFOSC2:
PESSTO_bias
PESSTO_calib

The night is not photometric

Weather conditions:
Start: moderate wind, high humidity
Middle and end: moderate wind

23:39 PESSTO_SkyFF_eve_UVi (just U because Vi were done in Nov17)
23:45 PESSTO_SkyFF_eve_BR

23:54 We start with EFOSC2. Image analysis
00:13 SN2016enp follow-up, gr13 with 1.5" slit, seeing ~1.9"
01:05 SN2016enp follow-up, gr13 with 1.5" slit, seeing ~1.4"-1.6"
01:52 Spectroscopic standard vma2, seeing ~1.3"-1.8"
02:16 SN2016iae follow-up, gr11+gr13 with 1.5" slit, seeing ~1.0"-1.3"(gr11),  ~1.0"-1.2"(gr16)
03:31 Switch to SOFI
03:45 bad SOFI configuration. 1h10m was lost in order to re-configure + Image analysis
04:55 SN2016hnk follow-up, JHK, seeing ~1.1"-1.3"
06:10 SN2016hvl follow-up, JHK, seeing ~1.37" (6:24=1.13", 6:36=1.18")
06:40 Switch to EFOSC2, Image analysis
06:46 SN2016hvl follow-up, gr11+gr16 with 1.0" slit, seeing ~ 1.0"
07:33 Spectroscopic standard LTT3864, seeing ~1.0"
07:59 OGLE16fht classification, seeing ~1.0"
08:33 End of the night





Comments